OpenStudy (alli6106):

A box contains 5 plain pencils and 5 pens. A second box contains 5 color pencils and 3 crayons. One item from each box is chosen at random. What is the probability that a plain pencil from the first box and a color pencil from the second box are selected? Write your answer as a fraction in simplest form.

OpenStudy (aum):

First box contains 5 plain pencils and 5 pens. If one item is picked at random, what is the probability that it is a plain pencil? There are a total of 10 items. 5 of them are plain pencils.

OpenStudy (alli6106):

it is in a fraction

OpenStudy (alli6106):

btw

OpenStudy (aum):

Can you answer my question?

OpenStudy (alli6106):

sure maybe if i can

OpenStudy (alli6106):

ummm hello

OpenStudy (aum):

I am waiting on your answer to the question in my first reply.

OpenStudy (aum):

Probability of an event happening = Number of ways the event can happen / Total number of outcomes First box contains 5 plain pencils and 5 pens. If one item is picked at random, what is the probability that it is a plain pencil? There are a total of 10 items. 5 of them are plain pencils. Therefore, Number of ways the event can happen = 5 and Total number of outcomes = 10 Probability of a plain pencil = 5/10 = 1/2 Similarly, probability of picking a color pencil from the second box = Number of color pencils / Total number of items in the second box = 5/8 The combined probability is: 1/2 * 5/8 = 5/16